OpenStudy (anonymous):

Evaluate Logarithmic. Show work. log0.0001

OpenStudy (anonymous):

hint": \(0.0001=10^{-4}\)

OpenStudy (anonymous):

log-10,000?

OpenStudy (anonymous):

no, try this \[\log_{10}(x)=y\iff 10^y=x\]

OpenStudy (anonymous):

on other words, \(\log(0.001)\) is the number \(y\) so that \(10^y=0.0001\)

OpenStudy (anonymous):

ok so whats next cause i did 10 by -4 lol

OpenStudy (anonymous):

it is not multiplication, it is the exponent if you want to know what \(\log(0.0001)\) is, you have to be able to solve \(10^y=0.0001\)

OpenStudy (anonymous):

since \(10^{-4}=0.0001\) you know \[\log(0.0001)=-4\]

OpenStudy (anonymous):

lol so its -4?

OpenStudy (anonymous):

yes, that is what it is

OpenStudy (anonymous):

thank you!